OpenStudy (anonymous):

Hi, i'm having to check my daughters work and make sure the answers are correct. But the math she is doing, i have never saw before in my life. So is there anyone on here that can help me?

OpenStudy (anonymous):

hii

OpenStudy (anonymous):

How are you today?

OpenStudy (anonymous):

lets see it

OpenStudy (anonymous):

Ok. Just a second.

OpenStudy (anonymous):

gud..yes bring it on :D

OpenStudy (anonymous):

Match the trig functions with the correct definitions: 1. Cos theta 2. Tan theta 3. Csc theta 4. Cot theta 5. Sec theta 6. Sin theta A. y/x B. y/r C. x/r D. r/y E. r/c AB. x/y

OpenStudy (anonymous):

Sorry, I can't help you with this, I don't know what this means

OpenStudy (anonymous):

I remember a little about having to do cos,tan,sin, etc but this theta stuff is new to me.

OpenStudy (anonymous):

i can help u

OpenStudy (anonymous):

Ok. Thank you so much.

OpenStudy (anonymous):

cos theta = x/r

OpenStudy (anonymous):

Same here, i only know the cos, tan, and sin, but thats it

OpenStudy (anonymous):

Me too. Ha. This new math curriculum is just crazy.

OpenStudy (anonymous):

I know! You never use it anyways! But good luck(:

OpenStudy (anonymous):

Thank you so much. Have a good day.

OpenStudy (anonymous):

You too!

OpenStudy (accessdenied):

 Theta (\(\theta\)) basically refers to the measure of some angle. Generally, we refer to the two sides that form the right angle as the 'legs of the right triangle', and the opposite side to the right angle the 'hypotenuse.' When we use the trigonometric functions with respect to the angle theta (in a right triangle), we look at where the sides are in comparison to theta. Sine refers to the ratio \(\large \frac{\text{opposite leg to } \theta}{\text{hypotenuse}}\) Cosine refers to the ratio \(\large \frac{\text{adjacent leg to } \theta}{\text{hypotenuse}}\) Tangent refers to the ratio \(\large \frac{\text{opposite leg}}{\text{adjacent leg}}\) Secant is just the reciprocal of Cosine (which sounds a little counter-intuitive) So, Secant refers to the ratio \(\large \frac{\text{hypotenuse}}{\text{adjacent leg}}\) Cosecant is the reciprocal of Sine. So, Cosecant refers to the ratio \(\large \frac{\text{hypotenuse}}{\text{opposite leg}}\) Cotangent is the reciprocal of Cosine. So, Cotangent refers to the ratio \(\large \frac{\text{adjacent leg}}{\text{opposite leg}}\) In terms of the x,y, and r, we generally refer to this setup:  (Analogous to an (x,y) graph) where the sides are basically the same as the previous diagram; x is adjacent to theta, y is opposite of theta, and r is the hypotenuse.